OpenStudy (anonymous):

find the distance from (2,-7) to the line defined buy y==2x-6. express as a radical or a number rounded to the nearest hundredth

OpenStudy (anonymous):

\[\sqrt((x-2)^2 + (y-(-7))^2) \] y = 2x-6 \[\sqrt((x-2)^2+(2x+1)^2)\] \[(x-2)^2 = x^2-4x+4; (2x+1)^2 = 4x^2+4x+1\] Those two added gives. \[5x^2+5\] So our distance formula is this. \[d=\sqrt(5x^2+5)\] Find the minimum distance by taking the derivative and setting it equal to 0. \[\frac{ 10x }{ 2\sqrt(5x^2+5) }=0\] x = 0 Plug x = 0 in the distance formula that we got. \[d=\sqrt(5)\]